OpenStudy (anonymous):

Its a physics question. Noone is answering in physics group so i am asking here. related to speed, velocity and accelertaion : A child drops a ball from a window. The ball strikes the ground in 3.0 seconds. What is the velocity of the ball, the instant before it hits the ground.

OpenStudy (irishboy123):

use \(v = u + at\) a = gravity

OpenStudy (irishboy123):

or just note that gravity = \(\frac{\Delta v}{\Delta t}\) = constant, so \(\Delta v = g.\Delta t\)